OpenStudy (anonymous):

\[17\sqrt[3]{2}-5\sqrt[3]{250}\]

OpenStudy (anonymous):

please show the work

OpenStudy (anonymous):

\[17\sqrt[3]{2} - 5\sqrt[3]{5^{2}*2}\] \[17\sqrt[3]{2}-25\sqrt[3]{2}\] \[-8\sqrt[3]{2}\]

OpenStudy (anonymous):

correct answer